How much would you have to raise the temperature of a copper wire(originally at 20 degrees) to increase its resistance by 12%?


1
Expert's answer
2020-11-24T05:42:39-0500

The ratio between the final and initial resistance must be


"\frac{R_f}{R_i}=1.12."

In terms of temperature coefficient of copper is will be


"\frac{R_f}{R_i}=\frac{R_{20}\cdot(1+\alpha T)}{R_{20}}=1.12,\\\space\\
T=\frac{1.12-1}{\alpha}=31°\text{C}."